OpenStudy (nudeoua):

The table shows the outputs y for different inputs x:

OpenStudy (jango_in_dtown):

@nudeoua hii

OpenStudy (jango_in_dtown):

Part A : Yes it is a function since all the x values have only one image each and neither of them have more than one images. Part B: we take the relation f(x)=2x+13 when x=7.f(x)=2*7+13=27 and the given function has the y value 8 when x=7 hence the relation f(x) has greater value when x=7 Part C: by the given condition 2x+13=75 so 2x=62 so x=31